# Inferring Change Points in High-Dimensional Regression via Approximate Message Passing

### v2

@@ Appendix A Proof of State Evolution
${\bm{\nu}}^{2t+1}=\lim_{N\to\infty}\frac{1}{N}\sqrt{\frac{\delta+1}{\delta}}{\mathop{\mathbb{E}}}\left[{\begin{bmatrix}{\bm{0}}_{L\times n}&\bm{B}^{\top}\end{bmatrix}}{\begin{bmatrix}{\bm{0}}_{n\times L}\\ f^{t}({\bm{Q}}^{2t}_{[-p]})\end{bmatrix}}\right]=\sqrt{\frac{\delta}{\delta+1}}\lim_{n\to\infty}\frac{1}{n}{\mathop{\mathbb{E}}}\left[\bm{B}^{\top}f^{t}({\bm{Q}}^{2t}_{[-p]})\right],$ (50)
@@
$=\lim_{n\to\infty}\frac{1}{n}{\mathop{\mathbb{E}}}\left[g^{t-1}\left({\bm{Q}}^{2t-1}_{[n]},q\left(\sqrt{\frac{\delta+1}{\delta}}\bm{Z}_{\bm{W}_{0},[n]},{\bm{\Psi}},\bm{\varepsilon}\right)\right)^{\top}g^{s-1}\left({\bm{Q}}^{2s-1}_{[n]},q\left(\sqrt{\frac{\delta+1}{\delta}}\bm{Z}_{\bm{W}_{0},[n]},{\bm{\Psi}},\bm{\varepsilon}\right)\right)\right]$ (54)
@@ Appendix B State Evolution Limits and Simplifications
B.1 State Evolution Dependence on Signal and Noise
In this section, we give a set of sufficient conditions for the existence of the limits in the state evolution equations (10)–(13), which also allow for removing the dependence of the state evolution on $\bm{B},\bm{\varepsilon}$. Assume (S0) on p.(S0), and also that the following assumptions hold:
@@ B.1 State Evolution Dependence on Signal and Noise
    As $n\to\infty$, the entries of the normalized change point vector ${\bm{\eta}}/n$ converge to constants $\alpha_{0},\dots,\alpha_{L}$ such that $10=\alpha_{0}<\alpha_{1}<\dots<\alpha_{L^{*}}=\ldots=\alpha_{L}=$ 1.
@@
The assumption (S1) is natural in the regime where the number of samples $n$ is proportional to $p$, and the number of degrees of freedom in the signals also grows linearly in $p$. Without change points, $f^{t},g^{t}$ can both be assumed separable without loss of optimality (due to (S0)). To handle the temporal dependence created by change points, we require $g^{t}_{i}$ to depend on $i$, for $i\in[n]$. However, Proposition 2 shows that it can be chosen to act row-wise, i.e., $g^{*t}_{i}(\bm{\Theta}^{t},{\bm{y}})=g^{*t}_{i}(\bm{\Theta}^{t}_{i},y_{i})$. This justifies (S2). When $g_{i}^{t}$ is chosen to be $g_{i}^{*t}$ for $i\in[n]$, the condition (S3) can be translated into regularity conditions on the prior marginals $\pi_{\bar{\Psi}_{i}}$ and distributional convergence conditions on the noise $\varepsilon_{i}$, as these are the only quantities that differ along the elements of the sets $\{\tilde{g}_{i}^{t}(\bm{Z}_{1},(\bm{V}^{t}_{\bm{\Theta}})_{1},\Psi_{\eta_{\ell}},\varepsilon_{i})\}_{i\in[\eta_{\ell},\eta_{\ell+1})}$ and $\{\partial_{1}{\tilde{g}_{i}^{t}(\bm{Z}_{1},(\bm{V}^{t}_{\bm{\Theta}})_{1},\Psi_{\eta_{\ell}},\varepsilon_{i})}\}_{i\in[\eta_{\ell},\eta_{\ell+1})}$ for $\ell\in[L^{*}-1]$. Under assumptions (S0)-(S3), the state evolution equations in (10)–(13) reduce to:
@@
where (S2) has allowed us to reduce the matrix products into sums of vector outer-products, and assumptions (S0)(S3) have removed the dependence on $\bm{B}$ and $\bm{\varepsilon}$ due to the law of large numbers argument in Lemma 4 of Bayati and Montanari (2011).
@@ B.2 Removing Dependencies on Noise on the RHS of (31) and (34)
Under assumptions (S0)(S3) above, the dependence of $\bm{V}_{\bm{\Theta}}^{t}$ on the RHS of (31)–(32) on $\bm{B},\bm{\varepsilon}$ can be removed. Moreover, the dependence of the RHS on $\bm{\varepsilon}$ through $\hat{{\bm{\eta}}}$ can be removed on a case-by-case basis. We expect this to hold, for example, under (S0)(S3) for estimators whose dependence on $\bm{\varepsilon}$ is via a normalized inner product. For example the estimator in (30) can be expressed as:
@@ C.1 Proof of Proposition 2
Lemma 6 (Extended Cauchy-Schwarz inequality, Lemma 2 in Lavergne (2008))
Let ${\bm{A}}\in{\mathbb{R}}^{n\times L}$ and ${\bm{B}}\in{\mathbb{R}}^{n\times L}$ be random matrices such that ${\mathop{\mathbb{E}}}\|{\bm{A}}\|_{F}^{2}<\infty,{\mathop{\mathbb{E}}}\|{\bm{B}}\|_{F}^{2}<\infty,$ and ${\mathop{\mathbb{E}}}[\bm{A}^{\top}\bm{A}]$ is non-singular, then
@@
Recall from (26) that the LHS of (70) is the objective we wish to minimise via optimal $f^{t}$. Indeed, by setting $f^{t}=f^{*t}$, (70) is satisfied with equality, which proves part 1 of Proposition 2.
@@
${\mathop{\mathbb{E}}}\left[\partial_{1i}\tilde{g}_{i}^{t}(\bm{Z}_{1},(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right]^{\top}$ (71)
=$\stackrel{{\scriptstyle(a)}}{{=}}{\mathop{\mathbb{E}}}_{(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}}\left\{{\mathop{\mathbb{E}}}_{\bm{Z}_{1},\bar{\varepsilon}}\left[\partial_{1i}\tilde{g}_{i}^{t}\left(\bm{Z}_{1},(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},\bar{\Psi}_{i},\bar{\varepsilon}\right)\bigg{|}(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}\right]\right\}^{\top}$
=$\stackrel{{\scriptstyle(b)}}{{=}}{\mathop{\mathbb{E}}}_{(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}}\left\{{\rm{{Cov}}}\left(\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}\right)^{-1}{\mathop{\mathbb{E}}}_{\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon}}\left[\left(\bm{Z}_{1}-{\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}]\right)\tilde{g}_{i}^{t}\left(\bm{Z}_{1},(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},\bar{\Psi}_{i},\bar{\varepsilon}\right)^{\top}\bigg{|}(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}\right]\right\}$
=$\stackrel{{\scriptstyle(c)}}{{=}}{\mathop{\mathbb{E}}}_{(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}}\left\{{\rm{{Cov}}}\left(\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}\right)^{-1}\right.$ (72)
$\hskip 28.45274pt\left.{\mathop{\mathbb{E}}}_{q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})}\left[{\mathop{\mathbb{E}}}_{\bm{Z}_{1}}\left[\left(\bm{Z}_{1}-{\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}]\right)g_{i}^{t}\left((\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right)^{\top}\bigg{|}(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right]\right]\right\}$
@@
$\hskip 28.45274pt\left.{\mathop{\mathbb{E}}}_{q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})}\left[\left({\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})]-{\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}]\right)g_{i}^{t}\left((\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right)^{\top}\right]\right\}$
=$\stackrel{{\scriptstyle(d)}}{{=}}{\mathop{\mathbb{E}}}\left\{\smash{\underbrace{{\rm{{Cov}}}\left(\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}\right)^{-1}\left({\mathop{\mathbb{E}}}\left[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right]-{\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}]\right)}_{=g_{i}^{*t}\left((\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right)}}\ g_{i}^{t}\left((\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right)^{\top}\right\}$
@@
$\lim_{n\to\infty}{\mathop{\mathbb{E}}}\left[\frac{1}{n}(g^{*t})^{\top}g^{*t}\right]-\underbrace{\lim_{n\to\infty}{\mathop{\mathbb{E}}}\left[\frac{1}{n}\left(g^{*t}\right)^{\top}g^{t}\right]}_{\left(\bar{{\bm{\nu}}}_{\bm{B}}^{t+1}\right)^{\top}}\underbrace{\left(\lim_{n\to\infty}{\mathop{\mathbb{E}}}\left[\frac{1}{n}(g^{t})^{\top}g^{t}\right]\right)^{-1}}_{\left(\bar{{\bm{\kappa}}}_{\bm{B}}^{t+1,t+1}\right)^{-1}}\underbrace{\lim_{n\to\infty}{\mathop{\mathbb{E}}}\left[\frac{1}{n}(g^{t})^{\top}g^{*t}\right]}_{\bar{{\bm{\nu}}}_{\bm{B}}^{t+1}}$ $\succcurlyeq 0.$ (76)
@@
$[(\bar{{\bm{\nu}}}_{\bm{B}}^{t+1})^{-1}]^{\top}\bar{{\bm{\kappa}}}_{\bm{B}}^{t+1,t+1}(\bar{{\bm{\nu}}}_{\bm{B}}^{t+1})^{-1}-{\bm{\Gamma}}_{\bm{B}}^{t+1}\succcurlyeq 0,$ (78)
@@ C.2 Computation of the Optimal Denoiser $g^{*t}$
Suppose $\bm{x}\in{\mathbb{R}}^{n}$ and $\bm{y}\in{\mathbb{R}}^{m}$ are jointly Gaussian with $\begin{bmatrix}\bm{x}\\ \bm{y}\end{bmatrix}\sim{\sf N}\left(\begin{bmatrix}\bm{\mu}_{x}\\ \bm{\mu}_{y}\end{bmatrix},\begin{bmatrix}{\bm{\Sigma}}_{x}&{\bm{\Sigma}}_{xy}\\ {\bm{\Sigma}}_{xy}^{\top}&{\bm{\Sigma}}_{y}\end{bmatrix}\right)$, then $({\bm{x}}\mid{\bm{y}}=\tilde{{\bm{y}}})\sim{\sf N}(\bm{\mu},{\bm{\Sigma}})$ where
@@
Let $\Gamma\sim\pi_{\Gamma}$ be a categorical variable supported on $[1,2,\dots,K]$. With $\bm{x}\in{\mathbb{R}}^{n}$ and $\bm{y}\in{\mathbb{R}}^{m}$, suppose

### v3

@@ Appendix A Proof of State Evolution
${\bm{\nu}}^{2t+1}=\lim_{N\to\infty}\frac{1}{N}\sqrt{\frac{\delta+1}{\delta}}{\mathop{\mathbb{E}}}\left[{\begin{bmatrix}{\bm{0}}_{L\times n}&\bm{B}^{\top}\end{bmatrix}}{\begin{bmatrix}{\bm{0}}_{n\times L}\\ f^{t}({\bm{Q}}^{2t}_{[-p]})\end{bmatrix}}\right],=\sqrt{\frac{\delta}{\delta+1}}\lim_{n\to\infty}\frac{1}{n}{\mathop{\mathbb{E}}}\left[\bm{B}^{\top}f^{t}({\bm{Q}}^{2t}_{[-p]})\right]$ (55)
@@
$=\lim_{n\to\infty}\frac{1}{n}{\mathop{\mathbb{E}}}\left[g^{t-1}\left({\bm{Q}}^{2t-1}_{[n]},q\left(\sqrt{\frac{\delta+1}{\delta}}\bm{Z}_{\bm{W}_{0},[n]},{\bm{\Psi}},\bm{\varepsilon}\right)\right)^{\top}\right$.
$\hskip 170.71652pt\left.\cdot\,g^{s-1}\left({\bm{Q}}^{2s-1}_{[n]},q\left(\sqrt{\frac{\delta+1}{\delta}}\bm{Z}_{\bm{W}_{0},[n]},{\bm{\Psi}},\bm{\varepsilon}\right)\right)\right]$ (59)
@@ Appendix B State Evolution Limits and Simplifications
B.1 Parametric Dependence of State Evolution on Signal and Noise
In this section, we give a set of sufficient conditions for the existence of the limits in the state evolution equations (10)–(13), which also allow for removing the parametric dependence of the state evolution on $\bm{B},\bm{\varepsilon}$. Assume (S0) on p.(S0), and also that the following assumptions hold:
@@ B.1 Parametric Dependence of State Evolution on Signal and Noise
    As $n\to\infty$, the entries of the normalized change point vector ${\bm{\eta}}/n$ converge to constants $\alpha_{0},\dots,\alpha_{L}$ such that $0=\alpha_{0}<\alpha_{1}<\dots<\alpha_{L^{*}}<\ldots<\alpha_{L}=1$.
@@
The assumption (S1) is natural in the regime where the number of samples $n$ is proportional to $p$, and the number of degrees of freedom in the signals also grows linearly in $p$. Without change points, $f^{t},g^{t}$ can both be assumed separable without loss of optimality (due to (S0)). To handle the heterogeneous dependence structure induced by change points, we require $g^{t}_{i}$ to depend on $i$, for $i\in[n]$. However, Proposition 2 shows that it can be chosen to act row-wise, that is, $g^{*t}_{i}(\bm{\Theta}^{t},{\bm{y}})=g^{*t}_{i}(\bm{\Theta}^{t}_{i},y_{i})$. This justifies (S2). When $g_{i}^{t}$ is chosen to be $g_{i}^{*t}$ for $i\in[n]$, the condition (S3) can be translated into regularity conditions on the prior marginals $\pi_{\bar{\Psi}_{i}}$ and distributional convergence conditions on the noise $\varepsilon_{i}$, as these are the only quantities that differ along the elements of the sets $\{\tilde{g}_{i}^{t}(\bm{Z}_{1},(\bm{V}^{t}_{\bm{\Theta}})_{1},\Psi_{\eta_{\ell}},\varepsilon_{i})\}_{i\in[\eta_{\ell},\eta_{\ell+1})}$ and $\{\partial_{1}{\tilde{g}_{i}^{t}(\bm{Z}_{1},(\bm{V}^{t}_{\bm{\Theta}})_{1},\Psi_{\eta_{\ell}},\varepsilon_{i})}\}_{i\in[\eta_{\ell},\eta_{\ell+1})}$ for $\ell\in[L^{*}-1]$. Under assumptions (S0)-(S3), the state evolution equations in (10)–(13) reduce to:
@@
where (S2) has allowed us to reduce the matrix products into sums of vector outer-products, and assumptions (S0)(S3) have removed the parametric dependence on $\bm{B}$ and $\bm{\varepsilon}$ due to the law of large numbers argument in Lemma 4 of Bayati and Montanari (2011).
@@ B.2 Removing Dependencies on Noise on the RHS of (31) and (34)
Under assumptions (S0)(S3) above, the parametric dependence of $\bm{V}_{\bm{\Theta}}^{t}$ on the RHS of (31)–(32) on $\bm{B},\bm{\varepsilon}$ can be removed. Moreover, the parametric dependence of the RHS on $\bm{\varepsilon}$ through $\hat{{\bm{\eta}}}$ can be removed on a case-by-case basis. We expect this to hold, for example, under (S0)(S3) for estimators whose parametric dependence on $\bm{\varepsilon}$ is via a normalized inner product. For example the estimator in (30) can be expressed as:
@@ C.1 Proof of Proposition 2
Lemma 6 (Matrix Cauchy-Schwarz inequality, Lemma 2 in Lavergne (2008))
Let ${\bm{A}},{\bm{B}}\in{\mathbb{R}}^{n\times L}$ be random matrices such that ${\mathop{\mathbb{E}}}\|{\bm{A}}\|_{F}^{2}<\infty,{\mathop{\mathbb{E}}}\|{\bm{B}}\|_{F}^{2}<\infty,$ and ${\mathop{\mathbb{E}}}[\bm{A}^{\top}\bm{A}]$ is non-singular, then
@@
Recall from (26) that the LHS of (75) is the objective we wish to minimise via optimal $f^{t}$. Indeed, setting $f^{t}=f^{*t}$, (75) is satisfied with equality, proving part 1 of Proposition 2.
@@
${\mathop{\mathbb{E}}}\left[\partial_{1i}\tilde{g}_{i}^{t}(\bm{Z}_{1},(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right]^{\top}$
$\stackrel{{\scriptstyle(a)}}{{=}}{\mathop{\mathbb{E}}}_{(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}}\left\{{\mathop{\mathbb{E}}}_{\bm{Z}_{1},\bar{\varepsilon}}\left[\partial_{1i}\tilde{g}_{i}^{t}\left(\bm{Z}_{1},(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},\bar{\Psi}_{i},\bar{\varepsilon}\right)\bigg{|}(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}\right]\right\}^{\top}$
$\stackrel{{\scriptstyle(b)}}{{=}}{\mathop{\mathbb{E}}}_{(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}}\left\{{\rm{{Cov}}}\left(\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}\right)^{-1}\right$.
$\hskip 14.22636pt\left.\cdot{\mathop{\mathbb{E}}}_{\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon}}\left[\left(\bm{Z}_{1}-{\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}]\right)\tilde{g}_{i}^{t}\left(\bm{Z}_{1},(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},\bar{\Psi}_{i},\bar{\varepsilon}\right)^{\top}\bigg{|}(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}\right]\right\}$
$\stackrel{{\scriptstyle(c)}}{{=}}{\mathop{\mathbb{E}}}_{(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}}\left\{{\rm{{Cov}}}\left(\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}\right)^{-1}\right$.
$\hskip 14.22636pt\left.\cdot{\mathop{\mathbb{E}}}_{q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})}\left[{\mathop{\mathbb{E}}}_{\bm{Z}_{1}}\left[\left(\bm{Z}_{1}-{\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}]\right)g_{i}^{t}\left((\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right)^{\top}\bigg{|}(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right]\right]\right\}$
@@
$\hskip 14.22636pt\left.\cdot{\mathop{\mathbb{E}}}_{q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})}\left[\left({\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})]-{\mathop{\mathbb{E}}}[\bm{Z}_{1}|(\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1}]\right)g_{i}^{t}\left((\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right)^{\top}\right]\right\}$
[…]
$\hskip 284.52756pt\left.\cdot g_{i}^{t}\left((\bar{\bm{V}}_{\bm{\Theta}}^{t})_{1},q(\bm{Z}_{1},\bar{\Psi}_{i},\bar{\varepsilon})\right)^{\top}\right\}$
@@
$\lim_{n\to\infty}{\mathop{\mathbb{E}}}\left[\frac{1}{n}(g^{*t})^{\top}g^{*t}\right]$
$\qquad-\underbrace{\lim_{n\to\infty}{\mathop{\mathbb{E}}}\left[\frac{1}{n}\left(g^{*t}\right)^{\top}g^{t}\right]}_{\left(\bar{{\bm{\nu}}}_{\bm{B}}^{t+1}\right)^{\top}}\underbrace{\left(\lim_{n\to\infty}{\mathop{\mathbb{E}}}\left[\frac{1}{n}(g^{t})^{\top}g^{t}\right]\right)^{-1}}_{\left(\bar{{\bm{\kappa}}}_{\bm{B}}^{t+1,t+1}\right)^{-1}}\underbrace{\lim_{n\to\infty}{\mathop{\mathbb{E}}}\left[\frac{1}{n}(g^{t})^{\top}g^{*t}\right]}_{\bar{{\bm{\nu}}}_{\bm{B}}^{t+1}}\succcurlyeq 0.$ (79)
@@
$(\bar{{\bm{\nu}}}_{\bm{B}}^{t+1})^{-1}]^{\top}\bar{{\bm{\kappa}}}_{\bm{B}}^{t+1,t+1}(\bar{{\bm{\nu}}}_{\bm{B}}^{t+1})^{-1}-{\bm{\Gamma}}_{\bm{B}}^{t+1}\succcurlyeq 0,$ (81)
@@ C.2 Computation of the Optimal Denoiser $g^{*t}$
Let $\bm{x}\in{\mathbb{R}}^{n}$ and $\bm{y}\in{\mathbb{R}}^{m}$ be jointly Gaussian, with
[…]
Then $({\bm{x}}\mid{\bm{y}}=\tilde{{\bm{y}}})\sim{\sf N}(\bm{\mu},{\bm{\Sigma}})$ where
@@
Let $\Gamma\sim\pi_{\Gamma}$ be a categorical variable supported on $1,2,\dots,K]$. With $\bm{x}\in{\mathbb{R}}^{n}$ and $\bm{y}\in{\mathbb{R}}^{m}$, suppose